An optical fibre is made of glass with a refractive index of 1.55 and is clad

with another glass with a refractive index of 1.51. The fibre has a core

diameter of 50µm and is used at a light wavelength of 0.8 µm.

(a) What NA does the fibre have?

(b) What is the acceptance angle?


1
Expert's answer
2021-01-13T11:34:55-0500

(a) By the definition of the numerical apperture, we have:


"NA=\sqrt{(n_1^2-n_2^2)}=\sqrt{(1.55^2-1.51^2)}=0.35."

(b) By the definition of the acceptance angle. we have:


"\theta_a=sin^{-1}(NA)=sin^{-1}(0.35)=20.5^{\circ}."

Answer:

(a) "NA=0.35."

(b) "\theta_a=20.5^{\circ}."